OpenStudy (anonymous):

Use the data in the following table, which summarizes results from 183 pedestrian deaths that were caused by accidents. If three different deaths are randomly selected without replacement, find the probability that they all involved intoxicated drivers. Pedestrian intoxicated? Yes No Driver intoxicated? Yes 20 32 No 80 51

OpenStudy (kropot72):

The total number of intoxicated drivers is 20 + 32 = 52 The probability of a driver being intoxicated is 52/183 The probability of there being no intoxicated drivers involved in the 3 deaths is \[P(zero\ intoxicated)=(1-\frac{52}{183})^{3}=0.3668\] \[P(3\ intoxicated)=1-0.3668=you\ can\ calculate\]

OpenStudy (anonymous):

Hmm. Okay thank you.

OpenStudy (kropot72):

You're welcome :)